The area of triangle ABC is 36 square units. If angle B = 30 degrees and C= 8 find the value of a
Answer by venugopalramana(3286)

AREA OF TRIANGLE =A=c*a*SIN(B)/2
36=8aSIN(30)=8a*1/2=4a
a=36/4=9